# Cross-Modal Temporal Fusion for Financial Market Forecasting

### v1

@@
Abstract.
Accurate financial market forecasting requires diverse data sources, including historical price trends, macroeconomic indicators, and financial news, each contributing unique predictive signals. However, existing methods often process these modalities independently or fail to effectively model their interactions. In this paper, we introduce Cross-Modal Temporal Fusion (CMTF), a novel transformer-based framework that integrates heterogeneous financial data to improve predictive accuracy. Our approach employs attention mechanisms to dynamically weight the contribution of different modalities, along with a specialized tensor interpretation module for feature extraction. To facilitate rapid model iteration in industry applications, we incorporate a mature auto-training scheme that streamlines optimization. When applied to real-world financial datasets, CMTF demonstrates improvements over baseline models in forecasting stock price movements and provides a scalable and effective solution for cross-modal integration in financial market prediction.
Multimodal learning, Large Language Models, Stock Market Forecasting, Model Interpretation
copyright: rightsretainedconference: PREPRINT; 2025; UKccs: Applied computing Forecastingccs: Information systems Expert systemsccs: Computing methodologies Search with partial observationsccs: Computing methodologies Continuous space searchccs: Computing methodologies Heuristic function constructionccs: Computing methodologies Natural language generationccs: Computing methodologies Neural networksccs: Computing methodologies Bayesian network modelsccs: Computing methodologies Modeling methodologies
1. Introduction
@@
The Efficient Market Hypothesis (EMH) (Fama, 1965, 1970) suggests that market efficiencies place limitations on the ability to consistently generate excess returns. In weak form efficiency, it is assumed that asset prices incorporate all information in past prices, making technical analysis ineffective; in semi-strong form, prices are assumed to reflect all public information, including historical prices, news, earnings reports, and economic data, therefore also rendering fundamental analysis ineffective; and in strong form efficiency prices are assumed to reflect both public and private information, making even insider trading ineffective. However, the EMH is controversial and there is ample empirical evidence of market inefficiencies (Barberis and Thaler, 2003), suggesting that it is possible to predict prices for excess returns.
The financial industry has been exploring prediction models since the early twentieth century (Cowles 3rd, 1933), continuously advancing these technologies through substantial financial investments. Traditional quantitative approaches mainly rely on historical time series data to forecast stock movements (Poon and Granger, 2003; Taylor, 2007). However, with the development of deep learning, more recent efforts have explored approaches to decompose complex market dynamics (Zhou et al., 2019; Pei et al., 2025) and capture stock interdependencies through attention mechanisms (Wang et al., 2019a; Cui et al., 2023).
Lately, advances in Natural Language Processing (NLP) have enabled the integration of unstructured textual data to enhance prediction models. For example, news (Hu et al., 2018; Picasso et al., 2019; Xu and Cohen, 2018) and social media content (Jin et al., 2020; Tolstikhin et al., 2021) can be analyzed for sentiment to generate a scoring matrix of positive-negative signals for each stock, which is then incorporated as a new input feature. These event-driven methods focus on extracting valuable patterns from event information for stock prediction.
Although prior work has achieved some success in stock prediction, three open challenges remain: (1) Heterogeneous data integration – existing methods (e.g., Poon and Granger, 2003; Zhou et al., 2019) tend to crudely aggregate multifrequency inputs (e.g., quarterly reports, daily price series, and real-time news) without aligning their temporal dependencies, which can lead to loss of signal or spurious correlations; (2) Superficial interpretability – existing use of attention mechanisms (Zhou et al., 2020; Wang et al., 2019b; Lin et al., 2021) fails to disentangle specific drivers (e.g., GDP trends vs. news) or provide actionable insights, making it difficult for practitioners to understand or trust predictions; (3) Inflexible training paradigms – characterized by rigid architectures with inefficient retraining and hyperparameter optimization (Zu et al., 2023; Choi et al., 2023), existing methods have limited ability to rapidly adapt to volatile market conditions.
@@
2. Literature Review
2.1. Multimodality
@@
Recent advances in multimodal learning have enabled the fusion of structured and unstructured data for financial forecasting. (Lim et al., 2021) introduced temporal fusion transformers to jointly model static covariates (e.g., sector metadata) and dynamic time series data, but their fixed temporal alignment struggles with low-frequency earnings reports. To address this, (Mandal et al., 2024) introduced a cross-modal transformer to align daily X (formerly Twitter) data with historical price trends, but their fixed temporal windows ignore intermodal frequency mismatches. A notable advancement is (Wood et al., 2021), which introduced the momentum transformer, which combines technical indicators with attention mechanisms to capture momentum-driven market regimes. However, like (Tavakoli et al., 2025), which fused numerical and textual datasets using cross-modal attention, these methods lack mechanisms to synchronize different granularity data with momentum shifts. Furthermore, despite their predictive performance, these models fail to provide clear explanations for their final results, limiting their interpretability and practical use in real financial decision-making.
2.2. Financial Time-series Forecast
@@
Modern financial time series forecasting increasingly leverages hybrid architectures. (Zhou et al., 2021) proposed Informer, a transformer variant optimized for time series prediction, which reduces the huge inference cost. For high-frequency trading data, such as the limit order book, (Kumar, 2023) designed a reinforcement learning framework with volatility-sensitive rewards. The efficiency of different model topologies has also been explored: some works, such as (Hou et al., 2021; Pei et al., 2025; You et al., 2024), use a graph-based topology, where stocks are represented as nodes with static and dynamic connections; in contrast, traditional multi-time series models follow a sequential topology, treating each stock as an independent time series without explicitly constructing their relationships. Recent work by (Wu et al., 2021) introduced Autoformer, which leverages autocorrelation to decompose market trends and seasonal effects; however, the model remains restricted to single-modal input and will not adapt to sudden market changes (e.g., a black swan event). Finally, most of these models directly use pre-processed data and do not address the complexities of processing multimodal unstructured data.
2.3. Interpretability
@@
For industrial applications, in particular, interpretability is a key requirement of financial forecasting models, as commercial vendors and regulators require actionable insights into model decisions. Post hoc explainability tools, such as attention maps (Lim et al., 2021) and saliency methods (Simonyan et al., 2014), have been introduced for deep learning models, but the explanations provided lack economic foundations. New deep learning architectures have also been introduced to improve interpretability. For example, to disentangle patterns, (Lin et al., 2021) proposed a Temporal Routing Adaptor with optimal transport, which learns distinct trading patterns and assigns stocks to patterns using dynamic routing. For multimodal settings, (Zhou et al., 2020) designed a Domain-Adaptive Neural Attention Network that aligns news sentiment trends with sector-specific price movements via cross-modal attention; however, although domain adversarial training improves robustness to distribution shifts, interpretability is reduced by masking attribution to specific modalities. Moreover, (Wang et al., 2019b) proposed a deep reinforcement learning framework that maps attention weights to a “buy-winners-sell-losers” trading strategy. By generating attention heatmaps from historical return patterns, their method is able to identify stocks with momentum behavior, but fails to disentangle causally significant factors (e.g., earnings surprises) from statistically incidental correlations.
@@
Figure 1. Overview of proposed CMTF. The framework integrates multimodal data (historical data, macro index, news, and financial reports). It employs Tensor Representation (extract tensor representation from unstructured data), Tensor Encoding (scale and preprocess the collected tensor), Tensor Interpretation (select important tensors), and a Transformer-based forecasting model (apply the optimal training scheme)
3. Preliminary
3.1. Notation
@@ 3.2. Task Definition
[…]
@@
[…]
@@ 4. Methodology
4.1. Tensor Representation
Figure 2. Tensor representation pipeline of financial reports and news articles.
@@ 4.1. Tensor Representation
To extract the tensor representation from unstructured news data, we implement a version of CatBoost’s gradient-boosting optimized decision trees (Prokhorenkova et al., 2018). The loss function shows as:
@@
$+\underbrace{\lambda||f_{m}||^{2}}_{\text{L2 Reg}}+\underbrace{\gamma(\mathbf{x}_{i})}_{\text{Encoding Stabilizer}}\Bigg{]}$
where $\mathbf{x}_{i}$ combines processed text signals and market technicals, including news sentiment, intraday price dynamics, and historical volatility; and $z^{n}_{i}$ represents the binary classification labels of next day price movements. Both $\mathbf{x}_{i}$ and $z^{n}_{i}$ contribute to the output tensor $Z^{n}$ from unstructured news data. $\theta$ denotes the parameters of the CatBoost model being optimized, while $F_{m-1}(\mathbf{x}_{i})$ is the ensemble prediction from the first $m{-}1$ trees, and $\gamma$ is for penalize.
@@
First, given an input document $U$, the LLM generates a five-dimensional rating vector $R\in\mathbb{R}^{5}$, where each dimension corresponds to a predefined semantic attribute. Formally,
(4) $R=f_{\text{rate}}(U)$
where $f_{\text{rate}}(\cdot)$ leverages contextual embeddings to infer rating scores.
Second, to ensure compatibility with downstream tasks, we map $R$ into a structured feature space optimized for predictive modeling. This transformation is defined as:
(5) $Z^{r}=f_{\text{proj}}(R)$
where $Z^{r}$ represents the final structured representation used for model training and inference.
After these steps, we have obtained four types of tensors for the next stage of encoding: tensors from historical data $Z^{h}$, macro index $Z^{m}$, news data $Z^{n}$, and financial reports $Z^{r}$.
4.2. Tensor Encoding
The influence of a specific event in an event-driven model will usually persist for an extended period, rather than being limited to a single point. Therefore, following the approaches of (Ortu et al., 2022; Saragih et al., 2024), we apply a weighted moving average (WMA) to model the decay of influence for data that do not have daily granularity. This assigns higher weights to more recent observations, allowing us to model the diminishing impact of an event over time:
(6) ${WMA}_{t}=\frac{\sum_{a=1}^{b}a\cdot S_{t-(b-a)}}{\sum_{a=1}^{b}a}$
Here, $b$ denotes the fixed window size; and $S_{t-(b-a)}$ represents the observation (e.g., news sentiment score) at time $t-a$. This approach allows us to model the extended influence over subsequent $a$ days.
@@ 4.2. Tensor Encoding
(7) $\text{Z}^{h,m,n,r}_{daily}=\text{TF}_{t}({Z}^{h,m,n,r})\quad$
@@ 4.3. Tensor Interpretation
(8) $\Phi_{\text{corr}}=\left\{d\in[1,D]\,\middle|\,\frac{1}{D-1}\sum_{\begin{subarray}{c}d^{\prime}=1\\ d^{\prime}\neq d\end{subarray}}^{D}|\rho(\mathbf{x}_{d},\mathbf{x}_{d^{\prime}})|<\tau_{\text{corr}}\right\}$
@@
(9) $\tilde{\mathcal{X}}_{t,d}=\begin{bmatrix}\mathcal{X}_{t^{{}^{\prime}},d}\\ \mathcal{X}_{t^{{}^{\prime}}-1,d}\end{bmatrix}\quad\forall d\in\Phi_{\text{corr}},\,t^{{}^{\prime}}\in\{2,\ldots,T^{{}^{\prime}}\}$
@@
We then solve the convex temporal group LASSO objective (Zhou et al., 2010; Wang et al., 2023):
(10) $\min_{W\in\mathbb{R}^{|\Phi_{\text{corr}}|\times N}}\underbrace{\frac{1}{2T^{{}^{\prime}}}\|Y-\tilde{\mathcal{X}}W\|_{F}^{2}}_{\text{Reconstruction Error}}+\alpha\underbrace{\sum_{d=1}^{|\Phi_{\text{corr}}|}\|W_{d}\|_{2}}_{\text{Cross-Target Sparsity}}$
@@
(11) $\Phi_{\text{final}}=\left\{d\in\Phi_{\text{corr}}\,\middle|\,\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\|W_{d}^{(k)}\|_{2}>0\right)\geq 0.8\right\}$
@@
Table 1. Data Introduction
@@
4.4. Event-Driven Forecast Model
@@
Let $H_{l}\in\mathbb{R}^{T\times d_{\text{model}}}$ denote the input to the attention layer at encoder layer $l$, where $d_{\text{model}}$ is the feature dimension of each token. This input $H_{l}$ includes the original feature embedding combined with positional encodings, following standard practice in transformer architectures (Vaswani et al., 2017; Devlin et al., 2019).
@@
(12) $Q_{h}=H_{l}W_{h}^{Q},\quad K_{h}=H_{l}W_{h}^{K},\quad V_{h}=H_{l}W_{h}^{V}$
(13) $\text{Attention}_{h}=\text{softmax}\left(\frac{Q_{h}K_{h}^{T}}{\sqrt{d_{k}}}\right)V_{h}$
(14) $\text{MHA}(H_{l})=\text{Concat}(\text{Attention}_{1},\ldots,\text{Attention}_{H})W^{O}$
@@
(15) $\text{FFN}(x)=\text{ReLU}(W_{1}x+b_{1})W_{2}+b_{2}$
@@
(16) $\text{PE}_{(pos,2\delta)}=\sin\left(\frac{pos}{10000^{2\delta/d_{\text{model}}}}\right)$
(17) $\text{PE}_{(pos,2\delta+1)}=\cos\left(\frac{pos}{10000^{2\delta/d_{\text{model}}}}\right)$
where $pos$ is the position of the token, and $\delta$ is the dimension index of the positional encoding.
@@
To enable rapid updates and efficient training, we use Optuna (Akiba et al., 2019) as the optimization framework. Optuna utilizes an asynchronous successive halving algorithm, which is equipped with two estimators to efficiently search for the local optimum in the hyperparameter space. The pruning criterion for each trial is defined as:
(19) $\text{Prune}(t)=\begin{cases}\text{True},&\text{if }\frac{r_{k}}{r_{k-1}}>\gamma^{1/\eta}\\ \text{False},&\text{otherwise}\end{cases}$
@@
Estimator 1: Tree-structured Parzen Estimator (TPE)
The TPE estimator models the probability of a set of hyperparameters $x$ given the value of the objective function $y$ as:
(20) $p(x|y)=\frac{\ell(x)}{\ell(x)+g(x)},\quad y\sim\text{Gamma}(k,\theta)$
where $\ell(x)$ and $g(x)$ represent the likelihood functions for good and bad hyperparameter configurations, respectively, and $y$ follows a Gamma distribution with shape $k$ and scale parameter $\theta$.
Estimator 2: Covariance Matrix Adaptation Evolution Strategy (CMA-ES)
The CMA-ES algorithm updates the mean $\mathcal{M}_{t+1}$ based on the weighted sum of the best candidates $y_{i:\lambda}$ from the population at time $t$ as:
(21) $\mathcal{M}_{t+1}=\mathcal{M}_{t}+\sigma_{t}\sum_{i=1}^{\mu}w_{i}y_{i:\lambda}$
where $\mathcal{M}_{t}$ is the mean at time $t$, $\sigma_{t}$ is the step size, $w_{i}$ are the weights of the candidates, and $y_{i:\lambda}$ denotes the best candidates in the population.
Table 2. Data Statistics
@@
5. Empirical Analysis of CMTF

### v2

@@
Abstract
Accurate forecasting in financial markets requires integrating diverse data sources, from historical prices to macroeconomic indicators and financial news. However, existing models often fail to align these modalities effectively, limiting their practical use. In this paper, we introduce a transformer-based deep learning framework, Cross-Modal Temporal Fusion (CMTF), that fuses structured and unstructured financial data for improved market prediction. The model incorporates a tensor interpretation module for feature selection and an auto-training pipeline for efficient hyperparameter tuning. Experimental results using FTSE 100 stock data demonstrate that CMTF achieves superior performance in price direction classification compared to classical and deep learning baselines. These findings suggest that our framework is an effective and scalable solution for real-world cross-modal financial forecasting tasks.
[…]
6025
1 Introduction
@@
The Efficient Market Hypothesis (EMH) [12, 13] suggests that market efficiencies place limitations on the ability to consistently generate excess returns. In weak form efficiency, it is assumed that asset prices incorporate all information in past prices, making technical analysis ineffective; in semi-strong form, prices are assumed to reflect all public information, including historical prices, news, earnings reports, and economic data, therefore also rendering fundamental analysis ineffective; and in strong form efficiency prices are assumed to reflect both public and private information, making even insider trading ineffective. However, the EMH is controversial and there is ample empirical evidence of market inefficiencies [2], suggesting that it is possible to predict prices for excess returns.
The financial industry has been exploring prediction models since the early twentieth century [7], continuously advancing these technologies through substantial financial investments. Traditional quantitative approaches mainly rely on historical time series data to forecast stock movements [25, 33]. However, with the development of deep learning, more recent efforts have explored approaches to decompose complex market dynamics [44, 23] and capture stock interdependencies through attention mechanisms [36, 8].
Lately, advances in Natural Language Processing (NLP) have enabled the integration of unstructured textual data to enhance prediction models. For example, news [16, 24, 41] and social media content [17, 34] can be analyzed for sentiment to generate a scoring matrix of positive-negative signals for each stock, which is then incorporated as a new input feature. These event-driven methods focus on extracting valuable patterns from event information for stock prediction.
Although prior work has achieved some success in stock prediction, three open challenges remain: (1) Heterogeneous data integration – existing methods [e.g., 25, 44] tend to crudely aggregate multifrequency inputs (e.g., quarterly reports, daily price series, and real-time news) without aligning their temporal dependencies, which can lead to loss of signal or spurious correlations; (2) Superficial interpretability – existing use of attention mechanisms [43, 20] fails to disentangle specific drivers (e.g., GDP trends vs. news) or provide actionable insights, making it difficult for practitioners to understand or trust predictions; (3) Inflexible training paradigms – characterized by rigid architectures with inefficient retraining and hyperparameter optimization [47, 6], existing methods have limited ability to rapidly adapt to volatile market conditions.
@@
The contributions and findings of this work have informed the development of Stratiphy’s emerging applications. Stratiphy is a wealth management platform for retail investors to build active portfolios using industry-leading trading strategies and risk management tools. [^1] Stratiphy can be found at https://www.stratiphy.io. The CMTF framework is being prototyped as an emerging application to develop new investment strategies for business and retail customers. These advances are essential to the continued success of Stratiphy and offer the social benefits of better financial investment and risk management for all. Code and data availability [^2] Code is available at https://github.com/PEIYUNHUA/CMTF; for data access, please contact Stratify..
2 Literature Review
2.1 Multimodality
@@
Recent advances in multimodal learning have enabled the fusion of structured and unstructured data for financial forecasting. [19] introduced temporal fusion transformers to jointly model static covariates (e.g., sector metadata) and dynamic time series data, but their fixed temporal alignment struggles with low-frequency earnings reports. To address this, [21] introduced a cross-modal transformer to align daily X (formerly Twitter) data with historical price trends, but their fixed temporal windows ignore intermodal frequency mismatches. A notable advancement is [39], which introduced the momentum transformer, which combines technical indicators with attention mechanisms to capture momentum-driven market regimes. However, like [32], which fused numerical and textual datasets using cross-modal attention, these methods lack mechanisms to synchronize different granularity data with momentum shifts. Furthermore, despite their predictive performance, these models fail to provide clear explanations for their final results, limiting their interpretability and practical use in real financial decision-making.
2.2 Financial Time-Series Forecasting
@@
Modern financial time series forecasting increasingly leverages hybrid architectures. [45] proposed Informer, a transformer variant optimized for time series prediction, which reduces the huge inference cost. For high-frequency trading data, such as the limit order book, [18] designed a reinforcement learning framework with volatility-sensitive rewards. The efficiency of different model topologies has also been explored: some works, such as [15, 42, 23], use a graph-based topology, where stocks are represented as nodes with static and dynamic connections; in contrast, traditional multi-time series models follow a sequential topology, treating each stock as an independent time series without explicitly constructing their relationships. Recent work by [40] introduced Autoformer, which leverages autocorrelation to decompose market trends and seasonal effects; however, the model remains restricted to single-modal input and will not adapt to sudden market changes (e.g., a black swan event). Finally, most of these models directly use pre-processed data and do not address the complexities of processing multimodal unstructured data.
2.3 Interpretability
@@
For industrial applications, in particular, interpretability is a key requirement of financial forecasting models, as commercial vendors and regulators require actionable insights into model decisions. Post hoc explainability tools, such as attention maps [19] and saliency methods [30], have been introduced for deep learning models, but the explanations provided lack economic foundations. New deep learning architectures have also been introduced to improve interpretability. For example, to disentangle patterns, [20] proposed a Temporal Routing Adaptor with optimal transport, which learns distinct trading patterns and assigns stocks to patterns using dynamic routing. For multimodal settings, [43] designed a Domain-Adaptive Neural Attention Network that aligns news sentiment trends with sector-specific price movements via cross-modal attention; however, although domain adversarial training improves robustness to distribution shifts, interpretability is reduced by masking attribution to specific modalities.
@@
Figure 1: Overview of proposed CMTF. The framework integrates multimodal data (historical data, macro index, news, and financial reports). It employs Tensor Representation (extract tensor representation from unstructured data), Tensor Encoding (scale and preprocess the collected tensor), Tensor Interpretation (select important tensors), and a Transformer-based forecasting model (apply the optimal training scheme).
3 Preliminary
3.1 Notation
@@ 3.2 Task Definition
$\text{\em true\_direction}^{i}_{t+1}=\begin{cases}1,&\text{if }p^{i}_{t+1}-p^{i}_{t}>0\\ 0,&\text{if }p^{i}_{t+1}-p^{i}_{t}<0\end{cases}$ (1)
@@
$\text{\em pred\_direction}^{i}_{t+1}=\begin{cases}1,&\text{if }\hat{p}^{i}_{t+1}-p^{i}_{t}>0\\ 0,&\text{if }\hat{p}^{i}_{t+1}-p^{i}_{t}<0\end{cases}$ (2)
@@ 4 Methodology
4.1 Tensor Representation
Figure 2: Tensor representation pipeline of financial reports and news.
@@ 4.1 Tensor Representation
To extract the tensor representation from unstructured news data, we implement a version of CatBoost’s gradient-boosting optimized decision trees [26]. The loss function shows as:
$\underset{\theta}{\arg\min}\sum_{m=1}^{M}\Bigg{[}\underbrace{\sum\ell(z^{n}_{i},F_{m-1}(\mathbf{x}_{i})+f_{m}(\mathbf{x}_{i}))}_{\text{Boosted Loss}}$ (3)
@@
where $\mathbf{x}_{i}$ combines processed text signals and market technicals, including news sentiment, intraday price dynamics, and historical volatility; and $z^{n}_{i}$ represents the binary classification labels of next day price movements. Both $\mathbf{x}_{i}$ and $z^{n}_{i}$ contribute to the output tensor $Z^{n}$ from unstructured news data. $\theta$ denotes the parameters of the CatBoost model being optimized, while $F_{m-1}(\mathbf{x}_{i})$ is the ensemble prediction from the first m1m{-}1 trees, and $\gamma$ is for penalize.
@@
First, given an input document $U$, the LLM generates a five-dimensional rating vector $R\in\mathbb{R}^{5}$. Second, to ensure compatibility with downstream tasks, we map $R$ into a structured feature space optimized for predictive modeling. This transformation is defined as:
$Z^{r}=f_{\text{proj}}(f_{\text{rate}}(U))$ (4)
where $f_{\text{rate}}(\cdot)$ leverages contextual embeddings to infer rating scores, and $Z^{r}$ represents the final structured representation used for model training and inference. Finally, we obtained four types of tensors for the next stage of encoding: tensors from historical data $Z^{h}$, macro index $Z^{m}$, news data $Z^{n}$, and financial reports $Z^{r}$.
4.2 Tensor Encoding
The influence of a specific event in an event-driven model will usually persist for an extended period, rather than being limited to a single point. Therefore, following the approaches of [22, 27], we apply a weighted moving average (WMA) to model the decay of influence for data that do not have daily granularity. This assigns higher weights to more recent observations, allowing us to model the diminishing impact of an event over time:
${WMA}_{t}=\frac{\sum_{a=1}^{b}a\cdot S_{t-(b-a)}}{\sum_{a=1}^{b}a}$ (5)
Here, $b$ denotes the fixed window size; and $S_{t-(b-a)}$ represents the observation (e.g., news sentiment score) at time tat-a. This approach allows us to model the extended influence over subsequent $a$ days.
@@ 4.2 Tensor Encoding
$\text{Z}^{h,m,n,r}_{daily}=\text{TF}_{t}({Z}^{h,m,n,r})\quad$ (6)
@@
$\Phi_{\text{corr}}=\left\{d\in[1,D]\,\middle|\,\frac{1}{D-1}\sum_{\begin{subarray}{c}d^{\prime}=1\\ d^{\prime}\neq d\end{subarray}}^{D}|\rho(\mathbf{x}_{d},\mathbf{x}_{d^{\prime}})|<\tau_{\text{corr}}\right\}$ (7)
@@
$\tilde{\mathcal{X}}_{t,d}=\begin{bmatrix}\mathcal{X}_{t^{{}^{\prime}},d}\\ \mathcal{X}_{t^{{}^{\prime}}-1,d}\end{bmatrix}\quad\forall d\in\Phi_{\text{corr}},\,t^{{}^{\prime}}\in\{2,\ldots,T^{{}^{\prime}}\}$ (8)
@@
We then solve the convex temporal group LASSO objective [37, 46]:
$\min_{W\in\mathbb{R}^{|\Phi_{\text{corr}}|\times N}}\underbrace{\frac{1}{2T^{{}^{\prime}}}\|Y-\tilde{\mathcal{X}}W\|_{F}^{2}}_{\text{Reconstruction Error}}+\alpha\underbrace{\sum_{d=1}^{|\Phi_{\text{corr}}|}\|W_{d}\|_{2}}_{\text{Cross-Target Sparsity}}$ (9)
@@
4.3 Tensor Interpretation
Figure 3: Demonstration of how tensor interpretation accumulates feature values over time. Here, we assume that CMTF undergoes monthly training iterations; a higher label count indicates greater importance for that period.
@@ 4.3 Tensor Interpretation
$\Phi_{\text{final}}=\left\{d\in\Phi_{\text{corr}}\,\middle|\,\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\|W_{d}^{(k)}\|_{2}>0\right)\geq 0.8\right\}$ (10)
@@
As demonstrated in Fig. 3, the relative frequency of each feature value determines its importance. As the training iterates monthly, it identifies key features that are important in specific time windows. This accumulation highlights stable or strongly correlated features, which improves the interpretation of cross-modal interactions over time.
Table 1: Data Summary.
@@
4.4 Event-Driven Forecast Model
@@
Let $H_{l}\in\mathbb{R}^{T\times d_{\text{model}}}$ denote the input to the attention layer at encoder layer $l$, where $d_{\text{model}}$ is the feature dimension of each token. This input $H_{l}$ includes the original feature embedding combined with positional encodings, following standard practice in transformer architectures [35, 10].
@@
$Q_{h}=H_{l}W_{h}^{Q},\quad K_{h}=H_{l}W_{h}^{K},\quad V_{h}=H_{l}W_{h}^{V}$ (11)
$\text{Attention}_{h}=\text{softmax}\left(\frac{Q_{h}K_{h}^{T}}{\sqrt{d_{k}}}\right)V_{h}$ (12)
$\text{MHA}(H_{l})=\text{Concat}(\text{Attention}_{1},\ldots,\text{Attention}_{H})W^{O}$ (13)
@@
$\text{FFN}(x)=\text{ReLU}(W_{1}x+b_{1})W_{2}+b_{2}$ (14)
@@
$\text{PE}_{(pos,2\delta)}=\sin\left(\frac{pos}{10000^{2\delta/d_{\text{model}}}}\right)$ (15)
$\text{PE}_{(pos,2\delta+1)}=\cos\left(\frac{pos}{10000^{2\delta/d_{\text{model}}}}\right)$ (16)
where pospos is the position of the token, and $\delta$ is the dimension index of the positional encoding.
@@
Figure 4: Hyperparameter search strategies over loss landscape: grid search, random search, and tree-structured parzen estimator (TPE).
@@
To enable rapid updates and efficient training, we use Optuna [1] as the optimization framework. Optuna utilizes an asynchronous successive halving algorithm, which is equipped with different estimators to search for the local optimum in the hyperparameter space. The pruning criterion for each trial is defined as:
$\text{Prune}(t)=\begin{cases}\text{True},&\text{if }\frac{r_{k}}{r_{k-1}}>\gamma^{1/\eta}\\ \text{False},&\text{otherwise}\end{cases}$ (18)
@@
Here, in the CMTF framework, we apply the default estimator, Tree-structured Parzen Estimator (TPE). The TPE estimator models the probability of a set of hyperparameters $x$ given the value of the objective function $y$ as:
$p(x|y)=\frac{\ell(x)}{\ell(x)+g(x)},\quad y\sim\text{Gamma}(k,\theta)$ (19)
where $\ell(x)$ and $g(x)$ represent the likelihood functions for good and bad hyperparameter configurations, respectively, and $y$ follows a Gamma distribution with shape $k$ and scale parameter $\theta$. As shown in Fig. 4, TPE demonstrates greater efficiency in identifying locally optimal hyperparameter combinations.
Table 2: Data Statistics.
@@
5 Empirical Analysis of CMTF